Determine the Concentration of H+
, OH-
, pH, and pOH of the following solutions:

1. 0.0100 M
2. 0.0200 M HOCl [ HOCl(aq) + H2O(l) ⇄ H3O+(aq) + OCl-(aq); Ka = 3.0 x 10 -8]
1
Expert's answer
2020-05-28T10:33:18-0400

HOCl is a weak acid that dissociates according to the following reaction:

HOCl(aq) + H2O(l) ⇄ H3O+(aq) + OCl-(aq).

If the equilibrium concentration of H3O+ is x, then the equilibrium constant Ka can be expressed in function of x:

"K_a = \frac{[H_3O^+][OCl^-]}{[HOCl]} = \frac{x^2}{c(HOCl)-x}" .

As Ka is very small, the can suggest that x<<c(HOCl). Then, the equilibrium concentration of [H3O+] (you can also use the [H+] notation) is "\sqrt{K_ac(HOCl)}" :

  1. c(HOCl)=0.0100 M : 1.73·10-5 M
  2. c(HOCl)=0.0200 M : 2.45·10-5 M.

Therefore, the pH of these solutions is:

  1. c(HOCl)=0.0100 M : pH = -log([H3O+]) = 4.76
  2. c(HOCl)=0.0200 M : pH = 4.61.

The pOH of water solutions can be found from the ionic product KW:


"pH + pOH = 14"


  1. c(HOCl)=0.0100 M : pOH = 14 - 4.76 = 9.24
  2. c(HOCl)=0.0200 M : pOH = 14 - 4.61 = 9.39.

Finally, the equilibrium concentration of OH- in these solutions:

  1. c(HOCl)=0.0100 M : [OH-] = 10-9.24 = 5.77·10-10 M
  2. c(HOCl)=0.0200 M : [OH-] = 10-9.39 = 4.08·10-10 M.

Answer:

  1. c(HOCl)=0.0100 M : [H+]=1.73·10-5 M, pH = 4.76, pOH = 9.24, [OH-] = 5.77·10-10 M
  2. c(HOCl)=0.0200 M : [H+]=2.45·10-5 M, pH = 4.61, pOH = 9.39, [OH-] = 4.08·10-10 M